# Magnetic Materials for Quantum Magnonics

### v1

@@ Abstract
Quantum magnonics studies the quantum properties of magnons — the quanta of spin waves — and their application in quantum information processing. Progress in this field depends on identifying magnetic materials with characteristics tailored to the diverse requirements of magnonics and quantum magnonics. For single-magnon excitation, its control, hybrid coupling, and entanglement, the most critical property is the ability to support long magnon lifetimes. This perspective reviews established and emerging magnetic materials — including ferromagnetic metals, Heusler compounds, antiferromagnets, altermagnets, organic and 2D van der Waals magnets, hexaferrites, and in particular yttrium iron garnet (YIG) — highlighting their key characteristics. YIG remains the benchmark, with bulk crystals supporting sub-microsecond Kittel-mode lifetimes and ultra-pure spheres achieving $\sim$18 µs for dipolar-exchange magnons at millikelvin temperatures. However, thin YIG films on gadolinium gallium garnet (GGG) substrates suffer from severe lifetime reduction due to substrate-induced losses. In contrast, YIG films on a new lattice matched, diamagnetic alternative, yttrium scandium gallium/aluminum garnet (YSGAG), overcomes these limitations and preserves low magnetic damping down to millikelvin temperatures. These advances provide a practical pathway toward ultralong-living magnons in thin films, enabling scalable quantum magnonics with coherent transport, strong magnon-photon, magnon-qubit coupling, and integrated quantum networks.
@@ I Introduction into quantum magnonics
Quantum computing promises transformative advances in computation, with potential impacts on cryptography, materials science, and artificial intelligence Nielsen and Chuang (2010); Arute et al. (2019). A central challenge for solid-state quantum technology is to identify an on-chip platform with a suitable nanoscale information carrier Chumak et al. (2022). Magnons — the bosonic quanta of spin-wave excitations in ordered magnets — are promising for this role. The emerging field of quantum magnonics Lachance-Quirion et al. (2019); Li et al. (2020, 2022); Chumak et al. (2022); Yuan et al. (2022) operates with quantum states of magnons, including single magnons, typically at millikelvin temperatures where thermal magnons are suppressed. Magnons cover gigahertz to terahertz frequencies Levchenko et al. (2024), scale to nanometer footprints Wang et al. (2019); Heinz et al. (2020), interface naturally with microwave circuitry Chumak, Serga, and Hillebrands (2017), and exhibit strongly nonlinear and nonreciprocal dynamics Chumak, Serga, and Hillebrands (2014); Wang et al. (2024); Pirro et al. (2021); Wang et al. (2023a); Zenbaa et al. (2025). These properties position magnons as promising candidates for quantum information processing, coherent transport, and entangled-state preparation Lachance-Quirion et al. (2019); Li et al. (2020, 2022); Chumak et al. (2022); Yuan et al. (2022).
@@
At room temperature (RT), macroscopic quantum phenomena in magnon gases are already accessible. Theory and experiment have established Bose–Einstein condensation of magnons Demokritov et al. (2006); Serga et al. (2014); Schneider et al. (2020, 2021), magnonic supercurrents Bozhko et al. (2016); Kreil et al. (2018), and the magnonic Aharonov–Casher effect Serha et al. (2023); Nakata et al. (2014). In antiferromagnets, quantum magnonics can reach the terahertz regime. Cavity coupled antiferromagnetic magnons have been demonstrated in collinear and canted phases Boventer et al. (2023), and terahertz cavity magnon polaritons have been realized in prototype platforms Kritzell et al. (2023). Long distance coherent antiferromagnetic spin-wave propagation at RT further supports device scale viability Wang et al. (2023b). These advances complement broader arguments for magnons as quantum excitations with frequency tunability and compatibility with hybrid architectures Yuan et al. (2022); Chumak et al. (2022); Li et al. (2020).
@@
Single-magnon operations in ferromagnetically ordered media require suppression of thermal magnon populations. At a frequency of 4 GHz, which is typical for superconducting qubits, high enough that three-magnon splitting is forbidden, and lifetimes are longer than at higher frequencies where viscous damping is strongerGurevich and Melkov (1996). To keep the thermal occupation below 0.01 by following the Bose–Einstein statistics, experiments must operate in dilution refrigerators at temperatures of about 40 mK or lowerClerk et al. (2010). The objective is to generate, control, and read out single magnons in hybrid systems that couple to photons, microwave cavities, phonons, or superconducting qubits Tabuchi et al. (2015); Wolski et al. (2020); Lachance-Quirion et al. (2020); Baity et al. (2021); Zuo et al. (2024).
@@
Property YIG Cherepanov, Kolokolov, and L’vov (1993); Serha et al. (2025a); Bertaut and Forrat (1956); Serga, Chumak, and Hillebrands (2010); Serha et al. (2025b); Geller and Gilleo (1957); Klingler et al. (2017, 2014); Maier-Flaig et al. (2017) Permalloy (Py)Demidov and Demokritov (2015); Kalarickal et al. (2006); Sebastian et al. (2015); Patton (1968) CoFeB Brunsch (1979); Liu et al. (2011); Conca et al. (2013); Zenbaa et al. (2025) Hematite Chen et al. (2025); Hamdi, Posva, and Grundler (2023); Wang et al. (2023b); Samuelsen and Shirane (1970) Hexaferrite BaM Song et al. (2010); Harris (2012); Malkinski (2012); Zhu et al. (2021); Popov et al. (2022) Heusler CMFSLiu et al. (2009); Trudel et al. (2010); Oogane et al. (2010); Sebastian et al. (2012) CrPS de Wal et al. (2023); Bud’ko et al. (2021); Lee et al. (2017); Freeman et al. (2025)
@@
Theory identifies roles for magnons as quantum interconnects and as probes of solid state quantum effects Kounalakis, Bauer, and Blanter (2022); Jiang et al. (2023); Dols et al. (2024); Dey et al. (2025). Further proposals predict steady state magnon squeezing and photon-magnon entanglement in cavity hybrids using Kerr nonlinearities and engineered couplings Yang et al. (2021); Toklikishvili et al. (2023); Sharma, Viola Kusminskiy, and Bittencourt (2024). Magnomechanical schemes show how squeezing can enhance correlations and light-matter control and enable nonclassical state preparation Lu et al. (2023); Zhang et al. (2024). Experimentally, key milestones include coherent coupling of fundamental magnon modes to superconducting resonators Lachance-Quirion et al. (2019); Li et al. (2022); Song et al. (2025a), qubit based quantum sensing Wolski et al. (2020), single magnon detection with a superconducting qubit Lachance-Quirion et al. (2020), full Wigner function tomography of a single magnon Xu et al. (2023), and NV based platforms that quantify magnon mediated NV-NV coupling Fukami et al. (2024); Borst et al. (2023). Although typical Kittel-mode magnon lifetimes Lachance-Quirion et al. (2020); Xu et al. (2023) of several hundred nanoseconds are shorter than the decoherence times of state-of-the-art superconducting qubits, which reach tens of microsecondsKrantz et al. (2019), they are sufficient for entanglement transfer and for interfacing heterogeneous quantum systems Tabuchi et al. (2015); Li et al. (2020); Yuan et al. (2022).
The most recent single-magnon experiments have been carried out on bulk spheres of yttrium iron garnet (YIG) Lachance-Quirion et al. (2020); Xu et al. (2023), which provides the longest coherence with Kittel-mode lifetimes generally below one microsecond Serga, Chumak, and Hillebrands (2010), underscoring its importance as the leading platform for quantum magnonics. Significant progress has been made in extending lifetimes at low temperatures, at which in ultra-pure bulk YIG, dipolar-exchange magnons reach about 18 µs at millikelvin temperatures Serha et al. (2025b).
Scaling from localized resonators to circuit-level links requires propagating magnons. Propagation enables spatial separation of single-magnon sources and detectors and allows the transport of quantum states across a chip Van Loo, Morris, and Karenowska (2018); Karenowska et al. (2015); Knauer et al. (2023). This shifts attention from bulk magnetic resonators to thin films, where dipolar-driven magnons can travel. However, thin films perform worse than bulk samples, as their higher surface-to-volume ratio amplifies the effect of surface defects. A further challenge in thin YIG films arises from the paramagnetism of GGG substrates at low temperatures Serha et al. (2025c, 2024); Guo et al. (2022); Jermain et al. (2017); Mihalceanu et al. (2018); Kosen et al. (2019). This limitation has recently been overcome by growing YIG films on a newly developed lattice-matched diamagnetic substrate Guguschev et al. (2025); Serha et al. (2025d). These advances enable on-chip platforms that excite and detect propagating magnons with planar resonators Wang et al. (2024). They also connect to ultralow-temperature studies of damping in bulk YIG Serha et al. (2025b); Kosen et al. (2019), to wavenumber-dependent damping in films Schmoll et al. (2025a), and to nanoscale propagation at cryogenic temperatures Knauer et al. (2023). Together, this progress establishes a foundation for quantum magnonic circuits that integrate superconducting qubits, photons, and phonons and exploit the nonlinear and nonreciprocal dynamics of magnons Barman et al. (2021); Flebus et al. (2024); Chumak et al. (2015, 2022).
II Available magnetic materials
Magnonics relies on ordered magnetic media that efficiently support spin waves. In this Perspective we briefly present and analyze different material platforms, including those that have been widely used for many years as well as more recent and highly attractive candidates. An overview is provided in Tab. 1, and representative examples are illustrated in Fig. 1.
@@ II Available magnetic materials
A common class of materials used to excite spin waves is ferromagnetic metals, valued for straightforward fabrication and high saturation magnetization (see Tab. 1). The most common examples are Permalloy ($\mathrm{Ni_{81}Fe_{19}}$)Demidov and Demokritov (2015); Kalarickal et al. (2006); Sebastian et al. (2015); Patton (1968), CoFeBTalmelli et al. (2020); Wojewoda et al. (2024); Brunsch (1979); Liu et al. (2011); Conca et al. (2013); Zenbaa et al. (2025) ($\mathrm{Co_{40}Fe_{40}B_{20}}$) and CoFeSchoen et al. (2016). However, their use in quantum magnonics is limited because free electrons provide a strong dissipation channel for spin waves, which leads to lifetimes of only a few nanosecondsKorenman and Prange (1972). Such short lifetimes pose a serious challenge for quantum operations with magnons, but they can still be of interest for studies in cryogenic magnonics Li et al. (2019). Working with lifetimes of this order requires strong coupling between systems to ensure sufficiently fast energy transfer.
Figure 1: (a) Temperature-dependent magnetization of hematite. Inset shows the easy-axis crystal structure for $T<T_{\mathrm{M}}$ (taken from Chen et al. (2025)). (b) Sketch of the antiferromagnetic (AFM) spin texture under microwave excitation. The magnetic field is applied along the $x$ direction. Inset shows a SEM image of the coplanar waveguide (CPW); scale bar $500~\mathrm{nm}$ (taken from Chen et al. (2025)). (c) Altermagnet: $d$-wave–type spin-split bands (upper left) and perovskite schematic with sublattice-dependent, anisotropic spin-dependent hoppings (red/blue arrows) (taken from Naka, Motome, and Seo (2025)). (d) Illustration of spin waves in a 2D van der Waals magnet. Magnon spintronics relies on manipulation and control of magnon spin transport from an injector to a detector. The electrical contacts for generating, controlling, and detecting magnons are shown as gold pads, while the van der Waals magnet is depicted by blue and yellow atoms (magnetic and non-magnetic, respectively). The spin is represented by a blue arrow whose orientation varies spatially to indicate the transported magnon (taken from Mañas-Valero et al. (2025)). (e) Atoms and spins in a bilayer of the van der Waals antiferromagnet CrPS4. Red and blue arrows indicate the local magnetic moments of the Cr atoms (turquoise). The interlayer (intralayer) exchange coupling is ferromagnetic (antiferromagnetic) (taken from de Wal et al. (2023)). (f) Lifetime $\tau$ of secondary dipolar–exchange magnons (DESW) at $f_{\mathrm{FMR}}/2$ as a function of temperature $T$ (logarithmic $T$-axis) for three YIG spheres. Illustration shows a schematic of three-magnon splitting — the $k=0$ uniform mode ($f_{\mathrm{FMR}}=3.17~\mathrm{GHz}$) splits into two counter-propagating magnons at $1.59~\mathrm{GHz}$ with $|k|\approx 3~\mathrm{rad}\,\mu\mathrm{m}^{-1}$ in the dipolar-exchange spin-wave regime (taken and adapted from Serha et al. (2025b)). (g) Experimental FMR linewidth $\Delta f$ as a function of the FMR frequency $f_{\mathrm{FMR}}$ for a $100~\mathrm{nm}$ YIG film on a GGG substrate. The half-solid points are measurement points up to $18~\mathrm{GHz}$ for temperatures down to $2~\mathrm{K}$ and up to $10~\mathrm{GHz}$ for temperatures below $2~\mathrm{K}$ (gray dashed lines mark these limits). The straight lines are Gilbert fits performed up to $18~\mathrm{GHz}$ and $10~\mathrm{GHz}$, respectively. Above these values, the linewidth shows non-Gilbert behavior versus $f_{\mathrm{FMR}}$. The linewidth broadening and deviation from Gilbert behavior are due to magnetization of the GGG substrate at low temperatures (inset), which creates an inhomogeneous stray field affecting the internal field of the YIG film (taken and adapted from Serha et al. (2025c)). (h) Crystal structure of diamagnetic YSGAG, an alternative substrate to paramagnetic GGG for YIG. The cubic lattice structure of YSGAG is similar to that of YIG, but with iron ions replaced at the octahedral sites by scandium, gallium, and aluminum, while tetrahedral sites are occupied by gallium and aluminum (taken from Serha et al. (2025d)).
@@
Heusler compounds, a class of mostly ternary or quaternary intermetallics with highly tunable electronic and magnetic properties Graf, Felser, and Parkin (2011), offer an improvement over conventional metallic ferromagnets. Among the most promising is CMFS $(\mathrm{Co_{2}Mn_{0.6}Fe_{0.4}Si})$. It exhibits longer spin-wave propagation lengths and lifetimes for dipolar spin waves,Tsymbal and Žutić (2019), yet these lifetimes remain orders of magnitude shorter than of the magnons currently used in quantum magnonic experiments. The finite electrical conductivity maintains a high spin-wave dissipation rate.

### v2

# Author to whom correspondence should be addressed: ]

@@ Abstract
Quantum magnonics studies the quantum properties of magnons — the quanta of spin waves — and their application in quantum information processing. Progress in this field depends on identifying magnetic materials with characteristics tailored to the diverse requirements of magnonics and quantum magnonics. For single-magnon excitation, its control, hybrid coupling, and entanglement, the most critical property is the ability to support long magnon lifetimes. This perspective reviews established and emerging magnetic materials — including ferromagnetic metals, Heusler compounds, antiferromagnets, altermagnets, organic and 2D van der Waals magnets, hexaferrites, europium chalcogenides, and in particular yttrium iron garnet (YIG) — highlighting their key characteristics. YIG remains the benchmark, with bulk crystals supporting sub-microsecond Kittel-mode lifetimes and ultra-pure spheres achieving $\sim$18 µs for dipolar-exchange magnons at millikelvin temperatures. However, thin YIG films on gadolinium gallium garnet (GGG) substrates suffer from severe lifetime reduction due to substrate-induced losses. In contrast, YIG films on a new lattice matched, diamagnetic alternative, yttrium scandium gallium/aluminum garnet (YSGAG), overcomes these limitations and preserves low magnetic damping down to millikelvin temperatures. These advances provide a practical pathway toward ultralong-living magnons in thin films, enabling scalable quantum magnonics with coherent transport, strong magnon-photon, magnon-qubit coupling, and integrated quantum networks.
@@ I Introduction into quantum magnonics
Quantum computing promises transformative advances in computation, with potential impacts on cryptography, materials science, and artificial intelligence Nielsen and Chuang (2010); Arute et al. (2019). A central challenge for solid-state quantum technology is to identify an on-chip platform with a suitable nanoscale information carrier Chumak et al. (2022). Magnons — the bosonic quanta of spin-wave excitations in ordered magnets — are promising for this role. The emerging field of quantum magnonics Lachance-Quirion et al. (2019); Li et al. (2020, 2022); Chumak et al. (2022); Yuan et al. (2022); Zhang (2023) operates with quantum states of magnons, including single magnons, typically at millikelvin temperatures, where thermal magnons are suppressed. Magnons cover gigahertz to terahertz frequencies Levchenko et al. (2026), scale to nanometer footprints Wang et al. (2019); Heinz et al. (2020), interface naturally with microwave circuitry Chumak et al. (2017), and exhibit strongly nonlinear and nonreciprocal dynamics Chumak et al. (2014); Wang et al. (2024); Pirro et al. (2021); Wang et al. (2023c); Zenbaa et al. (2025). These properties position magnons as promising candidates for quantum information processing, coherent transport, and entangled-state preparation Lachance-Quirion et al. (2019); Li et al. (2020, 2022); Chumak et al. (2022); Yuan et al. (2022).
@@
At room temperature (RT), macroscopic quantum phenomena in magnon gases are already accessible. Theory and experiment have established Bose–Einstein condensation of magnons Demokritov et al. (2006); Serga et al. (2014); Schneider et al. (2020, 2021), magnonic supercurrents Bozhko et al. (2016); Kreil et al. (2018), and the magnonic Aharonov–Casher effect Serha et al. (2023); Nakata et al. (2014). In antiferromagnets, quantum magnonics can reach the terahertz regime. Cavity-coupled antiferromagnetic magnons have been demonstrated in collinear and canted phases Boventer et al. (2023), and terahertz cavity magnon polaritons have been realized in prototype platforms Kritzell et al. (2023). Long-distance coherent antiferromagnetic spin-wave propagation at RT further supports device scaling viability Wang et al. (2023a). These advances complement broader arguments for magnons as quantum excitations with frequency tunability and compatibility with hybrid architectures Zhang (2023); Yuan et al. (2022); Chumak et al. (2022); Li et al. (2020).
@@
Single-magnon operations in ferromagnetically ordered media require suppression of thermal magnon populations. At a frequency of 4 GHz, which is typical for superconducting qubits and sufficiently high to forbid three-magnon splitting while providing longer lifetimes than at higher frequencies where viscous damping is stronger Gurevich and Melkov (1996), the thermal magnon occupation can be kept below 0.01 according to Bose–Einstein statistics only if experiments are performed in dilution refrigerators at temperatures of about 40 mK or lower Clerk et al. (2010). The objective is to generate, control, and read out single magnons in hybrid systems that couple to photons, microwave cavities, phonons, or superconducting qubits Tabuchi et al. (2015); Wolski et al. (2020); Lachance-Quirion et al. (2020); Baity et al. (2021); Zuo et al. (2024).
Theory identifies roles for magnons as quantum interconnects and as probes of solid state quantum effects Kounalakis et al. (2022); Jiang et al. (2023); Dols et al. (2024); Dey et al. (2025). Further proposals predict steady state magnon squeezing and photon-magnon entanglement in cavity hybrids using Kerr nonlinearities and engineered couplings Kostylev et al. (2019); Yang et al. (2021); Haghshenasfard and Cottam (2020); Toklikishvili et al. (2023); Sharma et al. (2024). Magnomechanical schemes show how squeezing can enhance correlations and light-matter control and enable nonclassical state preparation Lu et al. (2023); Zhang et al. (2024). Experimentally, key milestones include coherent coupling of fundamental magnon modes to superconducting resonators Lachance-Quirion et al. (2019); Li et al. (2022); Song et al. (2025b), qubit based quantum sensing Wolski et al. (2020), single magnon detection with a superconducting qubit Lachance-Quirion et al. (2020), full Wigner function tomography of a single magnon Xu et al. (2023), macroscopic Bell state between a spin system and a superconducting qubitXu et al. (2024), and NV-based platforms that quantify magnon mediated NV-NV coupling Fukami et al. (2024); Borst et al. (2023). Although the best Kittel-mode magnon lifetimes Lachance-Quirion et al. (2020); Xu et al. (2023) of several hundred nanoseconds are shorter than the decoherence times of state-of-the-art superconducting qubits, which reach tens of microsecondsKrantz et al. (2019), they are sufficient for entanglement transfer and for interfacing heterogeneous quantum systems Tabuchi et al. (2015); Li et al. (2020); Lachance-Quirion et al. (2019); Yuan et al. (2022); Zhang (2023). Recently, a qubit-free approach to quantum magnonics has been proposed, where Gaussian quantum states of ferromagnetic resonance magnons, including squeezedHaghshenasfard and Cottam (2016) and entangled states, are generated via parametric driving and magnonic nonlinearities, enabling quantum-state preparation and verification using conventional ferromagnetic resonance techniquesMycroft et al. (2026).
The most recent single-magnon experiments have been carried out on bulk spheres of yttrium iron garnet (YIG) Lachance-Quirion et al. (2020); Xu et al. (2023, 2024), which provides the longest coherence with Kittel-mode lifetimes generally below one microsecond Serga et al. (2010), underscoring its importance as the leading platform for quantum magnonics. Significant progress has been made in extending lifetimes at low temperatures, at which in ultra-pure bulk YIG, dipolar-exchange magnons reach about 18 µs at millikelvin temperatures Serha et al. (2025b).
A central challenge in integrating superconducting qubits with magnetic resonant systems is the strong sensitivity of conventional qubits to external magnetic fields, which significantly reduces their coherence times. Such fields are, however, required to tune magnetic resonators to microwave frequencies relevant for quantum applications. For instance, the Kittel mode of a YIG sphere at 4 GHz requires an external magnetic field of approximately 150 mT, a field strength that typical aluminum-based superconducting qubits cannot tolerate. Current experimental approaches therefore rely on spatially separating the magnetic resonator and the superconducting qubit, coupling them indirectly via a microwave cavity mode. This allows the YIG sphere to be biased by a magnetic field while keeping the qubit in a field-free environmentLachance-Quirion et al. (2020); Xu et al. (2023, 2024). This strategy could be further improved by employing planar superconducting resonators fabricated from materials that are more tolerant to magnetic fields such as niobium, niobium–tin or niobium-titanium, or by using qubits intrinsically robust against moderate magnetic fieldsPita-Vidal et al. (2020); Kringhøj et al. (2021); Günzler et al. (2025). An alternative route is the use of magnetic waveguides as magnonic media operating at or near zero external field. YIG waveguides have already demonstrated operation at frequencies around 1.7 GHzNikolaev et al. (2023). Upon cooling to cryogenic temperatures, the increase in saturation magnetizationSerha et al. (2025c) is expected to shift this frequency upward to approximately 3 GHz, placing it within a range that is well compatible with commonly engineered superconducting qubits.
Scaling from localized resonators to circuit-level links requires propagating magnons. Propagation enables spatial separation of single-magnon sources and detectors, and allows the transport of quantum states across a chip Van Loo et al. (2018); Karenowska et al. (2015); Knauer et al. (2023). This shifts attention from bulk magnetic resonators to thin films, where dipolar-driven magnons can travel. However, thin films perform worse than bulk samples, as their higher surface-to-volume ratio enhances the effect of surface defects. A further challenge in thin YIG films arises from the paramagnetism of GGG substrates at low temperatures Serha et al. (2025d, 2024); Guo et al. (2022); Jermain et al. (2017); Mihalceanu et al. (2018); Kosen et al. (2019). This limitation has recently been overcome by growing YIG films on a newly developed lattice-matched diamagnetic substrate Guguschev et al. (2025); Serha et al. (2025a). These advances enable on-chip platforms that excite and detect propagating magnons with planar resonators Wang et al. (2024). They also connect to ultralow-temperature studies of damping in bulk YIG Serha et al. (2025b); Kosen et al. (2019), to wavenumber-dependent damping in films Schmoll et al. (2025b), and to nanoscale propagation at cryogenic temperatures Knauer et al. (2023). Together, this progress establishes a foundation for quantum magnonic circuits that integrate superconducting qubits, photons, and phonons, and exploit the nonlinear and nonreciprocal dynamics of magnons Barman et al. (2021); Flebus et al. (2024); Chumak et al. (2015, 2022).
II Available magnetic materials
Magnonics relies on ordered magnetic media that efficiently support spin waves. In this Perspective, we briefly present and analyze different material platforms, including those that have been widely used for many years as well as more recent and highly attractive candidates. An overview is provided in Tab. 1, and representative examples are illustrated in Fig. 1. Further, the magnon lifetime in the discussed materials is commonly extracted from resonance measurements of the fundamental Kittel mode. The full width at half maximum (FWHM) $\Delta B$ of the magnetic resonance quantifies the magnetic losses of the system and allows one to determine the lifetime $\tau$ of the magnetic excitation via
$\tau=\frac{1}{\gamma\cdot\Delta B}\,$, (1)
where $\gamma$ is the gyromagnetic ratioGurevich and Melkov (1996). When comparing magnon damping and lifetimes across different magnetic materials, it is important to distinguish between the phenomenological Gilbert damping parameter $\alpha$ and the actual lifetime $\tau$ of a magnonic excitation. While $\alpha$ provides a convenient description of viscous damping in the Landau–Lifshitz–Gilbert framework and is commonly extracted from resonance linewidths, it does not uniquely determine magnon lifetimes in general. The magnetic linewidth $\Delta B$ of a system is related to the Gilbert damping parameter $\alpha$ by
$\Delta B(f)=\Delta B_{0}+\frac{4\pi\alpha}{\gamma}\,f\,$, (2)
where $\Delta B_{0}$ denotes the inhomogeneous linewidth broadening, corresponding to the linearly extrapolated linewidth at zero frequency ($f=0$).
In practice, linewidth broadening and magnon relaxation can arise from a variety of mechanisms beyond intrinsic viscous Gilbert damping, including impurity scattering, two-magnon scattering, magnon–phonon coupling, and slow-relaxor processes on rare-earth and $L$-state transition metal ion impurities, the latter being particularly relevant in YIG at cryogenic temperaturesGurevich and Melkov (1996); Pardavi-Horvath (2000). As a result, linewidth measurements may deviate substantially from those expected from a simple linear Gilbert damping picture, especially when comparing different magnon wavelengths, temperature regimes, or material classes.
@@ II Available magnetic materials
Property YIG Cherepanov et al. (1993); Serha et al. (2025c); Bertaut and Forrat (1956); Serga et al. (2010); Serha et al. (2025b); Geller and Gilleo (1957); Klingler et al. (2017, 2014); Maier-Flaig et al. (2017) Permalloy (Py)Demidov and Demokritov (2015); Kalarickal et al. (2006); Sebastian et al. (2015); Patton (1968) CoFeB Brunsch (1979); Liu et al. (2011); Conca et al. (2013); Zenbaa et al. (2025) Hematite Chen et al. (2025); Hamdi et al. (2023); Wang et al. (2023a); Samuelsen and Shirane (1970) Hexaferrite BaM Song et al. (2010); Harris (2012); Malkinski (2012); Zhu et al. (2021); Popov et al. (2022) Heusler CMFSLiu et al. (2009); Trudel et al. (2010); Oogane et al. (2010); Sebastian et al. (2012) CrPS de Wal et al. (2023); Bud’ko et al. (2021); Lee et al. (2017); Freeman et al. (2025)
@@
In antiferromagnets, the role of Gilbert damping and its relation to magnon lifetimes differ significantly from those in ferro- and ferrimagnets due to the two-sublattice nature of the magnetic orderGomonay et al. (2018). The dynamics are governed by the Néel order parameter and described by inertial equations of motion, with damping entering as a matrix rather than as a single scalar parameterKamra et al. (2018); Simensen et al. (2020). Consequently, direct comparisons of Gilbert damping parameters across different magnetic orders are generally not meaningful without accounting for the underlying relaxation mechanisms. Therefore, in this Perspective we focus on direct linewidth measurements of magnetic excitations in ferro- and ferrimagnets as a more reliable measure of the actual magnon lifetime.
@@
A common class of materials used to excite spin waves is ferromagnetic metals, valued for their straightforward fabrication, compatibility with standard nanofabrication techniques, and high saturation magnetization, which enables efficient spin-wave excitation and access to large frequency bandwidths (see Tab. 1). Prominent examples include Permalloy ($\mathrm{Ni_{81}Fe_{19}}$)Demidov and Demokritov (2015); Kalarickal et al. (2006); Patton (1968), CoFeB ($\mathrm{Co_{40}Fe_{40}B_{20}}$)Talmelli et al. (2020); Wojewoda et al. (2024); Brunsch (1979); Liu et al. (2011); Conca et al. (2013); Zenbaa et al. (2025), and CoFeSchoen et al. (2016). In particular, propagating spin-wave spectroscopy on epitaxial single-crystal Fe thin films revealed an effective Gilbert damping as low as $\alpha\approx 2.5\times 10^{-3}$, with lifetimes of up to 2 nsGladii et al. (2017). Even lower intrinsic damping has been demonstrated in carefully engineered CoFe alloys, where Gilbert damping values down to $\alpha\approx 5\times 10^{-4}$ were reportedSchoen et al. (2016). This reduction is attributed to a minimum in the electronic density of states at the Fermi level, which suppresses magnon–electron scattering and highlights the role of band-structure engineering in controlling dissipation in metallic systems. Their relatively large exchange stiffness and saturation magnetization further allow for short-wavelength magnons and high group velocities, advantageous for compact device architectures and broadband operationLevchenko et al. (2026).
At the same time, the presence of itinerant electrons in ferromagnetic metals introduces intrinsic dissipation channels through electron–magnon and electron–phonon scattering, resulting in magnetic damping that typically limits magnon lifetimes to the nanosecond regimeKorenman and Prange (1972). Despite short lifetimes ferromagnetic metals remain relevant for cryogenic magnonics and hybrid quantum platforms, particularly in regimes requiring fast dynamics, large saturation magnetization, electrical tunability, strong nonlinearities, or strong coupling to microwave photonsLi et al. (2019). Their metallic nature enables efficient electrical control via spin-transfer and spin–orbit torques and supports rapid energy exchange in strongly coupled hybrid systems, facilitating the integration of magnonic and electronic functionalities on a single chipZhang (2023).
[…]
@@
Heusler compounds, a class of mostly ternary or quaternary intermetallics with highly tunable electronic and magnetic propertiesGraf et al. (2011); Elphick et al. (2021); Mantion et al. (2024), offer an improvement over conventional metallic ferromagnets by combining high spin polarization, large saturation magnetization, and comparatively low magnetic damping. Among the most studied in the magnonics and spintronics communities are Co2-based full Heusler alloys, where half-metallic electronic structures and high crystalline order can suppress magnon dissipation channelsGuillemard et al. (2019).
A particularly promising example is the quaternary Heusler compound CMFS ($\mathrm{Co_{2}Mn_{0.6}Fe_{0.4}Si}$), for which low-damping spin-wave propagation and extended lifetimes have been demonstrated experimentallySebastian et al. (2012); Chumak (2019). Thin film studies of CFMS and related Co2FeGa0.5Ge0.5 (CFGG) alloys report effective Gilbert damping parameters in the range of $\alpha_{\rm eff}\sim 10^{-3}$ to $10^{-2}$, with evidence for both intrinsic and extrinsic damping contributions, such as two-magnon scatteringChumak et al. (2021). Additionally, substrate engineering has been shown to reduce Gilbert damping in CFAS (Co2FeAl0.5Si0.5) Heusler films, with optimal annealing conditions yielding $\alpha$ values significantly lower than in as-grown filmsLove et al. (2021).
Furthermore, theoretical and experimental studies on Co2MnZ (Z = Si, Ge, Sn, Sb) Heusler systems have identified pathways toward ultralow magnetic damping ($\alpha\sim 4\times 10^{-4}$$9\times 10^{-4}$), representing some of the lowest values reported for metallic ferromagnets and highlighting the crucial role of spin polarization and Fermi-level band structure in damping reductionGuillemard et al. (2019). More recently, ultra-low magnetic damping has been demonstrated in epitaxial Co2MnSi Heusler thin films, where transport measurements revealed a strong correlation between crystalline order, spin polarization, and reduced magnon dissipation, underscoring the potential of highly ordered Heusler alloys for low-loss magnon transport down to cryogenic temperaturesde Melo et al. (2021). In addition, other Co-based Heusler compounds, such as CoFe2Al, have also been reported to exhibit relatively low magnetic damping, suggesting a broader materials landscape for low-loss metallic magnonicsWang et al. (2023d).
Despite these improvements relative to conventional ferromagnetic metals, Heusler compounds still possess finite electrical conductivity, which continues to provide intrinsic dissipation channels via electron–magnon and electron–phonon scattering. Consequently, magnon lifetimes in optimized Heusler systems remain in the nanosecond regime. As a result, Heusler compounds represent an attractive and versatile materials platform for magnonics and spintronics, particularly in hybrid device architectures and electrically controlled magnon transport, while ongoing advances in materials optimization continue to expand their potential for long-lived quantum magnon applications.